OpenStudy (priyar):

Steel ruptures when a shear of 3.5 * 10 ^8 Nm^-2 is applied. The force neede to punch a 1cm diameter hole in a steel sheet 0.3cm thick is nearly?

OpenStudy (priyar):

@ganeshie8 @samigupta8 @parthkohli @mayankdevnani

OpenStudy (irishboy123):

isn't this just \(\tau = \dfrac{F}{A}\) ?

OpenStudy (priyar):

oh its that simple? shear means stress..? but i have learnt only categories in strain..not stress i am bit confused.. @IrishBoy123 ..

OpenStudy (irishboy123):

yes, the A is the cylindrical area that is being sheared, and that, if recall these things correctly, is that. bish bosh.

OpenStudy (priyar):

so would it be 3.5 * 10 ^8 * 3*10^-5 = 1.05 * 10^4N?

OpenStudy (irishboy123):

not too sure i follow all that but in letters \( F = \tau \times 2 \pi r t\) where t is metal thickness and r = radius of hole

OpenStudy (priyar):

i missed out pi in the abv exp. !

OpenStudy (priyar):

so F = 3.3 *10^4 N.. thats correct!

OpenStudy (priyar):

Thank You @IrishBoy123!!

OpenStudy (samigupta8):

BT shear is applied tangentially if we really want to punch the steel sheet then how can one do that without applying a normal force on it...?

OpenStudy (samigupta8):

@priyar do u know this?

OpenStudy (samigupta8):

Did you get my doubt?